# One-step Noisy Label Mitigation

### v1

@@ Abstract
Mitigating the detrimental effects of noisy labels on the training process has become increasingly critical, as obtaining entirely clean or human-annotated samples for large-scale pre-training tasks is often impractical. Nonetheless, existing noise mitigation methods often encounter limitations in practical applications due to their task-specific design, model dependency, and significant computational overhead. In this work, we exploit the properties of high-dimensional orthogonality to identify a robust and effective boundary in cone space for separating clean and noisy samples. Building on this, we propose One-step Anti-Noise (OSA), a model-agnostic noisy label mitigation paradigm that employs an estimator model and a scoring function to assess the noise level of input pairs through just one-step inference, a cost-efficient process. We empirically demonstrate the superiority of OSA, highlighting its enhanced training robustness, improved task transferability, ease of deployment, and reduced computational costs across various benchmarks, models, and tasks. Our code is released at https://github.com/leolee99/OSA.
@@ 1 Introduction
Noise mitigation aims to handle the detriment of noisy labels encountered during the training process. The advancement of large-scale pre-training has significantly increased data scale to the trillion level. Much of this data is sourced from the internet, inevitably introducing considerable noise, which severely impedes the training process. This poses a substantial challenge for robust model training in various tasks, such as cross-modal matching [1, 2], image-classification [3, 4], and image-retrieval [5].
Traditional noise mitigation approaches encounter several limitations that constrain their practical applicability: 1) Task specificity: Existing methods [1, 3, 6] are tailored to specific tasks, limiting their applicability across different tasks. 2) Model dependency: Most noise mitigation techniques [5, 7] are tightly coupled with specific models, requiring extensive modifications for adaptation to different models. 3) Computational cost: Numerous existing methods necessitate dual-model collaborations [1, 4] or multiple training passes [1], i.e., they require at least two backward passes per training step, effectively doubling the computational expense and substantially increasing the training burden (see Figure. 1a).
To tackle these challenges, we use an external estimator to assess the noise level of each sample, ensuring a model-agnostic approach. This estimator adjusts the training loss by reducing the influence of noisy samples, driving their weights toward zero. Furthermore, multimodal pre-trained models have demonstrated remarkable task transferability due to their strong semantic capabilities. For instance, CLIP [8] unifies the paradigms of image-text retrieval and image classification through a shared embedding space (see Figure. 1b). It converts category labels into sentences, maps them into the shared embedding space, and then calculates the cosine similarity with the image representation to perform image classification. Inspired by this, we leverage multimodal pre-trained models as estimators and apply the shared embedding space to enable task transfer. In this case, only one additional inference process is required for each sample, significantly reducing the computational overhead compared to performing an extra backward pass.
Nonetheless, this paradigm introduces a new challenge: how to accurately identify noise based solely on cosine similarity scores generated by estimators. An ideal solution is to find a decision boundary that separates clean samples from noise and accurately handles overlapping samples near the boundary. Existing methods [1, 9, 10, 2] typically attempt to build this boundary within the loss space, an isotropic space with uniform distribution, which creates only a narrow gap between noisy and clean samples. Moreover, the coarse handling of overlaps by integrating multi-model predictions often results in an unstable decision boundary. In contrast, the shared embedding space of pre-trained models is a high-dimensional, anisotropic space with an imbalanced distribution. Thus, a consideration is whether the properties of imbalanced anisotropic space can help to identify a more precise and robust decision boundary.
@@
In this work, we delve into the decision boundary of pre-trained models employed as estimators to accurately differentiate between clean and noisy samples. We first investigate the cosine similarity distributions of clean and noisy samples, calculated using the multimodal pre-trained models CLIP [8] and ALIGN [11], on two datasets with a 50% noise ratio: MS-COCO [12] and SDM, as shown in Figure. 1c-1f. SDM is a dataset with the images generated by the Stable Diffusion Model (SDM) [13] in some uncommon styles (see illustrations in Figure. 4). It is designed to explore how well-pre-trained models can distinguish unfamiliar domains that they rarely encounter during training. There are two interesting observations in Figure. 1c-1f: (1) The clean and noisy distributions of the same model on different datasets have a similar intersection point, suggesting the existence of a natural and stable boundary in distinguishing between clean and noisy samples. (2) The overlaps between the clean distribution and noisy distribution are minimal even in the unfamiliar domain dataset, indicating this boundary has strong potential for distinguishing between clean and noisy samples.
Building upon these two observations, we conduct an in-depth investigation and make the following contributions:
@@
1. We figure out the origin of the intersection, attributing it to the shift in the orthogonal boundary induced by the cone effect. Furthermore, we provide a theoretical framework that proves and elaborates the stability and precision of this boundary in separating noisy and clean samples.
2. We provide a detailed explanation of the reliability of pre-trained models in general noise recognition, even in unfamiliar domains, grounded in the analysis of the pre-training process.
3. Build on this, we introduce One-Step Anti-Noise (OSA), a general model-agnostic paradigm for noise recognition that requires only one-step inference. Specifically, we utilize a pre-trained model as the estimator to maintain a shared embedding space. A scoring function, designed based on the properties of high-dimensional orthogonality, is then used to accurately handle overlaps by directly assigning a learning weight to each sample’s loss according to its cosine similarity.
@@
4. We conduct comprehensive experiments across a variety of challenging benchmarks, models, and tasks, demonstrating the effectiveness, generalization capabilities, and efficiency of our method.
@@ 2 Boundary Principle Analysis
In Figure. 1c-1f, we observe a natural boundary emerging in the pre-trained model’s ability to distinguish between clean and noisy samples. In this section, we explain the principle of boundary forming from high-dimensional perspectives, and how robust it is in general noise mitigation.
2.1 Hypothesis: Interaction Boundary is Shifted from Orthogonal Boundary
We first elaborate on the gap extent between the positive and negative sides kept by the orthogonal boundary. Then, we present the reasoning behind the hypothesis that the intersection boundary in Figure. 1 is a shifted orthogonal boundary in the cone space.
The orthogonal boundary largely separates the positive and negative sides.
High-dimensional orthogonality is a general phenomenon caused by dimension disaster, where the angles between randomly selected vectors typically approximate 90 degrees, suggesting the cosine similarity that trends toward zero. For instance, in a 1024-dimensional space, the probability of two random vectors having a cosine similarity within $[-0.1,0.1]$ is approximately 99.86% (details are provided in Appendix. C.1). In this case, a natural boundary of cosine similarity zero forms, capably separating the positive side and negative side with a huge gap.
@@ 2.1 Hypothesis: Interaction Boundary is Shifted from Orthogonal Boundary
Recent literature [14, 15, 16] has demonstrated that the cone effect is a general phenomenon in deep neural networks, where the learned embedding subspace forms a narrow cone and the orthogonal boundary encounters a positive shift. Based on this, a hypothesis is that the interaction boundary in Figure. 1 is the shifted orthogonal boundary. To prove this, we simulate the process of selecting random vectors in high-dimensional space and randomly generate thousands of pairs mapped into the shared embedding space. We find that all similarity of these random vector pairs tends to a fixed value, with the low-variance cosine similarity almost lying in the middle of clean and noise distributions (see Table. 1). An interesting phenomenon is that if we compare the mean with the intersection points in Figure. 1c-1f, we find they are almost exactly the same. This suggests that the interaction boundary is highly likely to be a shifted orthogonal boundary in cone space.
2.2 Theoretical verification of the Interaction Boundary Origin
Here, we theoretically investigate whether the origin of the interaction boundary is a shifted orthogonal boundary. We first show that (i) contrastive learning separates clean and noisy samples on opposite sides of the orthogonal boundary and (ii) The relative relationships of pairs’ cosine similarity stays unchanged after transmitting into the narrow cone space. Based on (i) and (ii), we can confirm that the intersection boundary at the center of the clean and noisy distributions is the shifted orthogonal boundary.
For an initialized model intending to learn an embedding space, both clean and noisy samples are treated as orthogonal random vectors since lacking semantic perception ability in the initial space. During contrastive training process, given $N$ sample pairs $\{(x_{i},y_{i})\}^{N}_{i=1}$, the embedding space is optimized through the cross-entropy loss (Eq. 1).
To minimize $\mathcal{L}_{ce}$ during training, two subprocesses occur: the diagonal elements of the matrix (i.e., clean pairs) are optimized to the positive side of the orthogonal boundary, while the non-diagonal elements (equivalent to noise pairs) are optimized to the negative side. Consequently, the distributions of these two types of samples are on opposite sides of the orthogonal boundary.
Relative relationship unchanged in transmitting process.
@@ 2.2 Theoretical verification of the Interaction Boundary Origin
$\cos(F(u_{n}),F(v_{n}))<\cos(F(u),F(v))\approx\beta<\cos(F(u_{c}),F(v_{c}))\text{.}$ (2)
Theorem. 1 shows that the relative relationship of pairs in the original entire space, will not change after transmitting to the narrow cone space of the trained model, and there is always a boundary $\beta$ concentrated on most random vectors. In Appendix. C.2, we provide a detailed statement and proof of the Theorem.
2.3 Qualitative analysis of robustness and applicability
Next, we perform a qualitative analysis to explore (i) the robustness and generality of the boundary in distinguishing between clean and noisy samples, and (ii) how the boundary’s properties can be leveraged to achieve more reasonable and precise overlap handling.
@@ 2.3 Qualitative analysis of robustness and applicability
Although the boundary’s ability to distinguish clean and noisy samples is proven, its robustness and generality still require further exploration. For practical pre-training, it must maintain accuracy and robustness even in unfamiliar domain datasets. Since the capabilities of the pre-trained model are difficult to quantify, we conduct a qualitative analysis from the perspective of pre-trained model inference. The models pre-trained on millions of samples already possess somewhat semantic understanding capabilities. Given a positive pair from an unseen domain, due to the contrastive learning process during pre-training, it still has a strong likelihood of moving toward the positive side of the boundary, while the negative pair tends toward the negative side. Although the cosine similarity difference might be slight, as we have shown in Section. 2.1, the boundary constructs a significant gap from the perspective of high-dimensional orthogonality.
@@
Due to the properties of orthogonal boundary, as cosine similarity decreases and approaches zero from the positive side, the probability of positive samples sharply decreases. Therefore, we can design a scoring function to annotate the cleanliness of samples. This function should satisfy two requirements: for samples with cosine similarity less than or equal to zero, which are almost certainly noise, the function should assign them a weight of zero. For samples with cosine similarity greater than zero, the function gradient should increase rapidly as the cosine similarity moves further from zero.
@@ 3 Method
In this section, we present our One-Step Anti-Noise (OSA) paradigm with a workflow shown in Figure. 2. We first define the pair-based noise mitigation tasks for image-text matching, image classification, and image retrieval tasks in Sec. 3.1. Consequently, the detailed description of OSA is clarified in Sec. 3.2.
@@ 3.2 One-step Anti-Noise
The workflow of our noise mitigation approach OSA is depicted in Figure. 2. Initially, we utilize an estimator model to encode the input pair to a shared embedding space and continue to compute the cosine similarity between the paired embedding. Afterward, the cosine similarity is converted to a cleanliness score $w_{i},(0\leq w_{i}\leq 1)$ through a scoring function designed based on orthogonal properties (Section. 2.3). This score quantifies the clean degree of the sample, the smaller $w_{i}$ is, the noisier the sample.
@@
Figure 2: The workflow of OSA. In the anti-noise process, there are two phases: Scoring Phase and Training Phase. In the Scoring Phase, a pair is mapped to a shared embedding space by estimators. Then the cosine similarity is transformed to a weight $w$ by a scoring function. In the Training Phase, the weight $w$ is directly multiplied with the loss to instruct the optimization.
@@ 3.2.1 Estimator Model
In our approach, the Estimator Model must satisfy two critical requirements: 1) effectively mapping input pairs into a unified embedding space and 2) possessing basic semantic understanding capabilities. To meet these requirements, we employ CLIP [8], a commonly used multimodal pre-trained models, as our estimator model. It is equipped with a text encoder $\phi_{t}(\cdot)$ and an image encoder $\phi_{v}(\cdot)$, enabling it to perform basic zero-shot tasks efficiently.
@@
While we have performed a qualitative analysis of the zero-shot pre-trained model’s robustness on out-of-domain data in Section. 2.3, and shown strong robustness for edge cases in Figure. 1, considering the domain diversity in real-world scenarios, we provide an optional Domain Adaptation (DA) approach to enhance the estimator model’s adaptability when encountering edge domains. Following NPC [2], we first employ a Gaussian Mixture Model (GMM) coupled with strict selection thresholds to ensure the absolute cleanliness of the chosen samples. We afterward implement a warm-up phase with few steps, allowing the estimator model to better understand the semantics of the target domain. Notably, this trick is only optional for our methods. Through multiple experiments, we found that even without domain adaptation, the zero-shot CLIP model performs exceptionally well across various scenarios.

### v2

# Debiased Orthogonal Boundary-Driven Efficient Noise Mitigation

@@ Abstract
Mitigating the detrimental effects of noisy labels on the training process has become increasingly critical, as obtaining entirely clean or human-annotated samples for large-scale pre-training tasks is often impractical. Nonetheless, existing noise mitigation methods often encounter limitations in practical applications due to their task-specific design, model dependency, and significant computational overhead. In this work, we exploit the properties of high-dimensional orthogonality to identify a robust and effective boundary in cone space for separating clean and noisy samples. Building on this, we propose One-Step Anti-noise (OSA), a model-agnostic noisy label mitigation paradigm that employs an estimator model and a scoring function to assess the noise level of input pairs through just one-step inference. We empirically validate the superiority of OSA, demonstrating its enhanced training robustness, improved task transferability, streamlined deployment, and reduced computational overhead across diverse benchmarks, models, and tasks. Our code is released at https://github.com/leolee99/OSA
Debiased Orthogonal Boundary-Driven Efficient Noise Mitigation
Hao Li1thanks: These authors contribute equally to this work., Jiayang Gu2, Jingkuan Song3thanks: Correspondence., An Zhang4, Lianli Gao5 1Washington University in St. Louis, 2University of Warwick, 3Tongji University, 4University of Science and Technology of China, 5University of Electronic Science and Technology of China {18th.leolee, jiayang.barrygu, jingkuan.song}@gmail.com
@@ 1 Introduction
Increasing the scale of data is a relatively efficient way to improve model performance in many domains Kaplan et al. (2020). As existing works tend to harvest data from the internet by query matching Sharma et al. (2018), structural attribute matching Schuhmann et al. (2022), it inevitably introduces label noise into the training process. The incompatibility between the data and the label hinders positive samples to learn correct embeddings, thus wasting the distributional diversity brought by large-scale data and the cost of collecting them. This poses a substantial challenge for robust model training in various tasks, such as cross-modal matching Huang et al. (2021); Zhang et al. (2024), image classification Yu et al. (2019), and image retrieval Liu et al. (2021).
Noise mitigation methods aim to eliminate the detriment effects from noisy samples during model training. However, existing noise mitigation approaches encounter several limitations that constrain their practical applicability: 1) Task specificity: Existing methods Huang et al. (2021); Sun et al. (2021); Ibrahimi et al. (2022) are tailored to specific tasks, limiting their applicability across different tasks. 2) Model dependency: Most noise mitigation techniques Liu et al. (2021); Yang et al. (2023a) are tightly coupled with specific models, requiring extensive modifications for adaptation to different models. 3) Computational cost: Numerous existing methods necessitate dual-model collaborations Huang et al. (2021); Yu et al. (2019) or multiple training passes Huang et al. (2021), i.e., they require at least two backward passes per training step, effectively doubling the computational expense and substantially increasing the training burden (see Figure. 1a). Motivated by the remarkable generalization capabilities of multimodal pre-trained models such as CLIP Radford et al. (2021), several studies Feng et al. (2024); Wei et al. (2024); Zhang et al. (2024); Liang et al. (2023) start leveraging these pre-trained models for noise mitigation. However, these approaches still suffer from aforementioned limitations, including task specificity Feng et al. (2024); Wei et al. (2024); Liang et al. (2023), model dependency Wei et al. (2024), and excessive computational demands Zhang et al. (2024), making them hard to utilize in practical scenarios.
@@
To tackle these challenges, we use an external estimator to assess the noise level of each sample, ensuring the target model-agnostic. This estimator reduces the influence of noisy samples by reducing their weights of training loss closer to zero. We leverage multimodal pre-trained models as the estimator due to their revealed strong semantic capabilities and task transferability. For instance, CLIP Radford et al. (2021) unifies the paradigms of image-text retrieval and image classification through a shared embedding space (see Figure. 1b). It converts category labels into sentences and then calculates the cosine similarity with the image representation to perform image classification. In this case, only one additional inference process is required for each sample, significantly reducing the computational overhead compared to performing an extra backward pass.
Nonetheless, this paradigm introduces a new challenge: how to accurately identify noise based solely on cosine similarity scores inferred by estimators. An ideal solution is to find a decision boundary in cosine space that can separate clean and noisy samples. Existing methods (Huang et al., 2021; Li et al., 2020; Feng et al., 2024; Zhang et al., 2024) typically attempt to build this boundary within the loss space, an isotropic space with uniform distribution, which creates only a narrow gap between noisy and clean samples. More critically, the coarse handling of overlaps by integrating multi-model predictions often results in an unstable decision boundary. In contrast, the shared embedding space of pre-trained models is a high-dimensional space, and its corresponding cosine similarity space is an anisotropic space with the non-uniform distribution. Thus, a consideration is whether the properties of imbalanced anisotropic space can help to identify a more precise and robust decision boundary.
In this work, we delve into the issue of decision boundary selection in anisotropic cosine spaces for pre-trained models to be efficient noise estimators. Theoretically, a cosine similarity of zero—i.e., an orthogonal boundary—should serve as a natural decision threshold in anisotropic cosine spaces to separate clean and noisy samples. To validate this hypothesis, we empirically analyze the cosine similarity distributions of clean and noisy samples using multimodal pre-trained models CLIP Radford et al. (2021) and ALIGN Jia et al. (2021) across two datasets: MSCOCO and SDM (Stable Diffusion Model Rombach et al. (2022)), both with a 50% noise ratio. The SDM dataset, comprising images generated in uncommon artistic styles (see Figure. 4), is designed to test the robustness of pre-trained models in distinguishing noisy samples from unseen domains.
Surprisingly, as shown in Figure 1c-1f, the empirically optimal decision boundary deviates significantly from the theoretical orthogonal threshold zero, limiting its usage in practical applications. Despite this deviation, there are also two interesting observations: (1) the intersection points of clean and noisy distributions remain consistent for the same model across different datasets, suggesting the existence of a stable, dataset-irrelevant boundary. (2) even on the SDM dataset—where models encounter unfamiliar domains—the overlap between clean and noisy distributions remains minimal, indicating the boundary’s robustness in distinguishing noisy samples.
Building on these two observations, we aim to reveal the underlying mechanisms and provide a theoretical guidance for fully exploiting the potential of pre-trained models in noise mitigation. Our key contributions are as follows:
@@
1. We theoretically figure out the origin of the intersection, attributing it to the shift of orthogonal boundaries induced by the cone effect and elaborate the stability and accuracy of this boundary in separating noisy and clean samples.
2. Building on our findings, we develop One-Step Anti-noise (OSA), an efficient and model-agnostic paradigm for noise recognition that requires just one-step inference.
@@
3. We conduct comprehensive experiments across a variety of challenging benchmarks, models, and tasks, demonstrating the effectiveness, generalization capabilities, and efficiency of our findings and introduced methods.
@@ 2 Boundary Principle Analysis
In Figure 1c-1f, we observe a natural boundary emerging in the pre-trained model’s ability to distinguish between clean and noisy samples. In this section, we explain the principle of boundary forming from high-dimensional perspectives, and how robust it is in general noise mitigation.
2.1 Hypothesis: Intersection Boundary is Shifted from Orthogonal Boundary
We first analyze the margin between the positive and negative regions defined by the orthogonal decision boundary. We then provide the rationale for our hypothesis that the intersection boundary in Figure 1 can be interpreted as a shifted orthogonal boundary within the conic space.
The orthogonal boundary clearly partitions the positive and negative regions.
High-dimensional orthogonality is a well-known phenomenon often referred to as the curse of dimensionality, where the angles between randomly selected vectors tend to concentrate around $90^{\circ}$, implying cosine similarities close to zero. For example, in a 1024-dimensional space, the probability that two random vectors have a cosine similarity within [0.1,0.1][-0.1,0.1] is approximately 99.86% (see Appendix C.1). This concentration effect induces a natural boundary at zero cosine similarity, which separates vectors with positive and negative similarity values with a wide margin.
@@ 2.1 Hypothesis: Intersection Boundary is Shifted from Orthogonal Boundary
Recent literature Liang et al. (2022a); Bogolin et al. (2022); Ethayarajh (2019) has demonstrated that the cone effect is a general phenomenon in deep neural networks, where the learned embedding subspace forms a narrow cone and the orthogonal boundary encounters a positive shift. Based on this, a hypothesis is that the intersection boundary in Figure 1 is the shifted orthogonal boundary. To prove this, we simulate the process of selecting random vectors in high-dimensional space and randomly generate thousands of pairs mapped into the shared embedding space. We find that all similarity of these random vector pairs tends to a fixed value, with the low-variance cosine similarity almost lying in the middle of clean and noise distributions (see Table. 1). An interesting phenomenon is that if we compare the mean with the intersection points in Figure 1c-1f, we find they are almost identical, suggesting that the intersection is highly likely to be a shifted orthogonal boundary in cone space.
2.2 Theoretical Verification of Intersection Origin
Here, we theoretically examine whether the intersection boundary originates from a shifted orthogonal boundary. Specifically, we demonstrate that (i) the relative ordering of pairwise cosine similarities remains invariant under projection into the conic space, and (ii) contrastive learning separates clean and noisy samples into opposite half-spaces defined by the orthogonal boundary. From (i) and (ii), it follows that the intersection boundary between the noise-free and noisy distributions corresponds to the shifted orthogonal boundary.
Relative relationship unchanged in propagation process.
@@ 2.2 Theoretical Verification of Intersection Origin
$\cos(F(u_{n}),F(v_{n}))$ $<\cos(F(u),F(v))$ (1)
$\quad\approx\beta<\cos(F(u_{c}),F(v_{c})).$
Theorem. 1 shows the pairs’ relative relationship in the original entire space remain unchanged after propagating to the narrow cone space of the trained model, and there is always a boundary $\beta$ concentrated on most random vectors. Appendix. C.2 provides a detailed statement and proof.
For an initialized model to learn an embedding space, both clean and noisy samples are treated as orthogonal random vectors since lacking semantic perception ability in the initial space. During contrastive training process, given $N$ sample pairs $\{(x_{i},y_{i})\}^{N}_{i=1}$, the embedding space is optimized through the cross-entropy loss:
To minimize $\mathcal{L}_{ce}$ during training, two subprocesses occur: the diagonal elements of the matrix (i.e., clean pairs) are optimized to the positive side of the orthogonal boundary, while the non-diagonal elements (equivalent to noise pairs) are optimized to the negative side. Consequently, the distributions of these two types of samples are on opposite sides of the orthogonal boundary. Combining with the Theorem. 1, we can determine that the interaction boundary between clean and noise in post-trained model is the shifted of orthogonal boundary.
2.3 Discussion on Applicability
@@
To further explore the boundary’s generalization ability, especially fine-tuning in unfamiliar domains, we conduct a qualitative analysis on how our boundary can still be effective in this more practical scenario. Given a positive pair from an unseen domain, due to the contrastive learning process during pre-training, it still has a strong likelihood of moving toward the positive side of the boundary, while the negative pair tends toward the negative side. Although the cosine similarity difference might be slight, as we have shown in Section. 2.1, the boundary constructs a significant gap from the perspective of high-dimensional orthogonality.
@@
Since orthogonal boundary properties, as cosine similarity decreases and approaches zero from the positive side, the probability of positive samples sharply decreases. Therefore, we can design a scoring function to annotate the cleanliness of samples. This function should satisfy two requirements: for samples with cosine similarity less than or equal to zero, which are almost certainly noise, the function should assign them a weight of zero. For samples with cosine similarity greater than zero, the function gradient should increase rapidly as the cosine similarity moves further from zero.
@@ 3 Method
In this section, we present our One-Step Anti-noise (OSA) paradigm with a workflow shown in Figure. 2. We first define the pair-based noise mitigation tasks in Sec. 3.1. Afterward, we clarify OSA in Sec. 3.2.
@@ 3.2 One-step Anti-Noise
The workflow of our noise mitigation approach OSA is depicted in Figure. 2. Initially, we utilize an estimator model to encode the input pair to a shared embedding space and continue to compute the cosine similarity between the paired embeddings. Afterward, the cosine similarity is converted to a cleanliness score $w_{i},(0\leq w_{i}\leq 1)$ through a scoring function designed based on orthogonal properties (Section. 2.3). This score quantifies the clean degree of the sample, the smaller $w_{i}$ is, the noisier the sample.
@@
Figure 2: OSA workflow. In the scoring phase, the estimator model calculates the semantic similarity score for input data pairs. Then the score is calibrated by debiasing the effect from the shifted orthogonal boundary. By passing through a scoring function, the in-batch loss is reweighted by the weight $w$.
@@ 3.2.1 Estimator Model
In our approach, the Estimator Model must satisfy two critical requirements: 1) effectively mapping input pairs into a unified embedding space and 2) possessing basic semantic understanding capabilities. To meet these requirements, we employ CLIP (Radford et al., 2021), a commonly used multimodal pre-trained models, as our estimator model. It is equipped with a text encoder $\phi_{t}(\cdot)$ and an image encoder $\phi_{v}(\cdot)$, enabling it to perform basic zero-shot tasks efficiently.
@@
While we have performed a qualitative analysis of the zero-shot pre-trained model’s robustness on out-of-domain data in Section. 2.3, and shown strong robustness for edge cases in Figure. 1, considering the domain diversity in real-world scenarios, we provide an optional Domain Adaptation (DA) approach to enhance the estimator model’s adaptability when encountering edge domains. Following NPC (Zhang et al., 2024), we first employ a Gaussian Mixture Model (GMM) coupled with strict selection thresholds to ensure the absolute cleanliness of the chosen samples. We afterward implement a warm-up phase with few steps, allowing the estimator model to better understand the semantics of the target domain. Notably, this trick is only optional for our methods. Through multiple experiments, we found that even without domain adaptation, the zero-shot CLIP model performs exceptionally well across various scenarios.